OpenStudy (anonymous):

How does a^(5/3)*(3/2)=^(5/2)?

OpenStudy (amistre64):

hard to tell

OpenStudy (lalaly):

(5/3)*(3/2)= 15/6 =5/2 :S

OpenStudy (anonymous):

Thanks lalaly lol... I can't believe i missed that...

OpenStudy (amistre64):

as far as being able to decipher that, kudos lol

OpenStudy (amistre64):

must be one of those mathematical notation classes i missed ;)